# Identification of Craving Maps among Marijuana Users via the Analysis of Functional Brain Networks with High-Order Attention Graph Neural Networks

### v4

@@ 2.2.3 Brain network reconstruction and community detection
Functional connectivity assesses brain connections between various regions. To enhance network connectivity and community structure, we utilize the final output embedding, denoted by $\textbf{H}^{c}$, for network reconstruction. This is accomplished through an inner product operation, as depicted in Eq. (14):
$\textbf{B}=\sigma(\textbf{H}^{c}\cdot\textbf{H}^{c^{T}}),$ (14)
where $\textbf{B}\in\mathbb{R}^{n\times n}$ represents the reconstructed matrix for ROIs. We utilize $\sigma(\cdot)$, for a nonlinear transformation to construct a weighted reconstructed matrix. Subsequently, for each subject’s connectivity matrix, denoted by $\textbf{A}^{(i)}_{r}(t)$ (where $i$ represents the subject and $r$ denotes the loop iteration), we perform element-wise multiplication with the weighted reconstructed matrix, $\textbf{B}_{r}$. This computation is shown in Eq. (15) and results in the final average weighted functional connectivity matrix.
$\textbf{W}^{FC}=\frac{1}{N}\sum_{t=1}^{N}\textbf{A}^{(i)}_{r}(t)\odot B_{t},$ (15)
After obtaining the average weighted functional connectivity $\textbf{W}^{FC}$, we further analyze the brain community structure for community detection, which $\odot$ denotes the element-wise product. The concept of community detection in graphs is based on the analysis of edge density. The Girvan-Newman (GN) modularity algorithm is utilized to ascertain whether based on the idea that there is a higher degree of similarity between nodes within a community and a lower degree of similarity between nodes outside the community [31, 32]. The GN algorithm follows these steps: (1) calculate the edge betweenness for each edge; (2) remove the edge with the highest edge betweenness; (3) recalculate the edge betweenness for the remaining edges in the network; (4) repeat steps (3) and (4) until any vertex in the network becomes a community. We can compute modularity $M$ and define community $C$ in the graph as:
$M=\frac{1}{2m}\sum_{ij}(W^{FC}_{ij}-P_{ij})\delta(C_{i},C_{j}).$ (16)
where $m$ represents the total number of edges in the graph, and $B$ denotes the adjacency matrix. We compute the expected number of edges $P_{ij}=p=2m/\left[n(n-1)\right]$ between $i$ and $j$ in Eq. (16). The $\delta$-function, considering vertices $i$ and $j$ belonging to the same community ($C_{i}=C_{j}$), aids in clustering within the community structure and assessing modularity.
2.2.4 Objective function
@@
We can employ the output embedding in Eq. (5) and Eq. (13) throughout the training phase. Eq. (17) and Eq. (18) to execute a binary classification task, incorporating the final non-linear transformation $\sigma(\cdot)$ as follows:
$\hat{\textbf{Y}}_{c}=\sigma(\textbf{W}_{c}^{T}\textbf{H}_{i}^{c}+b_{c}),$ (17)
$\hat{\textbf{Y}}_{MixGAT}=\sigma(\textbf{W}_{a}^{T}\textbf{H}_{i}^{a}+b_{a}),$ (18)
@@
where $\textbf{W}{c}$ and $\textbf{W}{a}$ represent the trained parameters, and $b_{c}$ and $b_{a}$ denotes biases. Subsequently, our objective is to minimize the overall loss function derived from the losses of two sub-models. The first sub-model pertains to the MixHop attention loss, denoted as $\mathcal{L}_{MixGAT}$, while the second sub-model involves the combination fusion model loss from Eq. (19) output embedding, denoted as $\mathcal{L}_{c}$. The combined loss can be expressed as:
@@
where Y represents the ground truth label, while $\hat{\textbf{Y}}_{MixGAT}$ and $\hat{\textbf{Y}}_{c}$ represent the class predictions for marijuana users and healthy controls, respectively. By combining two $\mathcal{L}_{MixGAT}$ and $\mathcal{L}_{c}$, we can better leverage the fusion information in the total loss $\mathcal{L}$ not just sub-model loss.
@@ 3 Results
3.1 Dataset
  • Marijuana-323: In this study, we extended the prior work of Kaustubh R. et al  [33]. The research involved the collection of data from two datasets of magnetic resonance imaging (MRI) images comprising 125 and 198 participants, recruited either from the non-medical marijuana-seeking community or from hospitalized patients. Before and after the MRI scans, participants completed the marijuana craving questionnaire [34], marijuana withdrawal checklist, and marijuana problem investigation  [35]. The data underwent reprocessing from 3T imaging fMRI data, and an average time series of 90 ROIs was calculated for each subject based on the Stanford altas  [36]. In our final training and testing, there were 195 long-term marijuana users and 128 healthy control individuals.
  • HCP dataset: To assess the effectiveness of HOGANN, we evaluated an external dataset sourced from the human connectome project (HCP) S1200 [37], comprising 1096 rs-fMRI scans of young adults. The S1200 constitutes the major releases from 2012 to 2015 from Washington University in St. Louis, Missouri, United States. All subjects were recruited and approved by the Institutional Review Board (IRB) review committee at Washington University in St. Louis. Detailed information regarding HCP data preprocessing can be accessed from the project’s website(https://www.humanconnectome.org/). The HCP dataset encompasses 598 individuals who have used marijuana, with an average age of 28.76 ± 3.69, and 493 individuals who have never used marijuana, with an average age of 28.82 ± 3.72. This data preparation utilized each subject’s first session and excluded 5 rs-fMRIs containing fewer than 1200 frames. The cortical surfaces of each subject were partitioned into 22 brain regions serving as ROIs  [38], and blood oxygenation level-dependent (BOLD) time series signals were normalized using z-scores  [19].
3.2 Baseline methods
We conducted comparisons among linear methods, such as logistic regression (LR) and SVM [39], and graph-based methods with some state-of-the-art GNNs, and graph-based methods, including GNNs, including GCN, GraphSAGE [40], GAT, GraphSAGE, and BrainGNN [21] based on brain network construction pipelines. Linear classifiers were evaluated using both L1 and L2 regularization terms consistent with the original paper.
  • LR with penalty is a linear classification method that uses Lasso regularization to prevent overfitting. The penalty in L1 or L2 regularization is determined by the sum of the absolute values or the sum of the absolute values or the sum of squares of the feature weights. This technique drives the weights of specific features toward zero, thereby promoting feature selection.
  • SVM with penalty is a supervised learning algorithm that will be used for classification and regression analysis tasks. The combination of L1 and L2 penalty terms is referred to as Elastic Net. This is achieved by adding a linear combination of L1 and L2 penalty terms and adjusting the corresponding mixing parameter.
  • GCN is a semi-supervised graph convolutional network model designed to learn node representations by leveraging neighboring information. It is commonly applied to learning graph structural data.
  • GAT extends GNNs by introducing an attention mechanism that enables dynamic adjustment of node relationship weights during the learning process.
  • GraphSAGE is an inductive graph convolutional network capable of predicting unseen nodes. By employing a sampling strategy, the model learns node embeddings based on neighboring nodes’ features rather than relying on the entire graph.
  • BrainGB  [20] BrainGB is a platform founded on GNN methods. It facilitates predictions across functional and structural neuroimaging modalities. GNNs are employed to model relationships between different brain regions
Table 1: Overall classification performance
Dataset Model AUC (%) Accuracy (%) Precision (%) Recall (%)
L1 LR $77.7$ $77.2$ $78.0$ $77.7$
L2 LR $75.4$ $72.8$ $75.2$ $75.4$
L1 SVM $79.2$ $77.9$ $79.1$ $79.2$
L2 SVM $77.6$ $75.6$ $77.5$ $77.7$
Marijuana-323 GCN $77.7$ $71.5$ $71.9$ $71.5$
GAT $81.0$ $73.8$ $74.0$ $73.9$
GraphSAGE $77.3$ $74.6$ $75.4$ $74.6$
BrainGB $83.4$ $75.0$ $73.6$ $69.3$
HOGANN (w/o MixHop) $81.9$ $76.8$ $76.8$ $76.9$
HOGANN (w/o LSTM) $85.1$ $76.2$ $76.1$ $76.2$
HOGANN 85.1 80.7 81.0 80.7
L1 LR $62.8$ $62.7$ $63.3$ $62.7$
L2 LR $64.9$ $64.8$ $65.2$ $64.8$
L1 SVM $64.6$ $64.7$ $64.7$ $64.7$
L2 SVM $69.6$ $60.9$ $64.5$ $60.9$
HCP GCN $49.9$ $50.5$ $25.5$ $50.5$
GAT $65.6$ $68.2$ $68.3$ $68.2$
GraphSAGE $54.4$ $53.9$ $55.7$ $53.9$
BrainGB $54.4$ $54.5$ $52.4$ $51.9$
HOGANN (w/o MixHop) $70.4$ $64.8$ $66.0$ $64.8$
HOGANN (w/o LSTM) $66.3$ $67.2$ $67.3$ $67.2$
HOGANN 75.5 71.2 71.7 71.7
3.2.1 Linear classifier
We compared the proposed linear classifier methods, including LR and SVM, with L1 and L2 regularization. Table 1 illustrates the model performance across different approaches in the marijuana-323 dataset. Our experiments reveal that among the linear classification methods—LR with L1 regularization, LR with L2 regularization, and SVMs with both L1 and L2—the SVM with L1 regularization achieved the highest classification accuracy at 79.2%. However, the same linear methods did not yield superior results in the HCP dataset. Only SVM with L2 regularization achieved a classification accuracy of 69.6%. These classification results across the two datasets indicate the presence of nonlinear structures among brain regions in the fMRI scans of long-term cannabis users, which linear classifiers fail to discern based on functional connectivity features.
3.2.2 Graph-level classification
We compared the proposed method with other graph-based methods such as GCN, GAT, GraphSAGE, and BrainGB as the baseline models. Table 1 shows that GAT and BrainGB achieve a better area under the curve (AUC) scores than linear classifier methods in the marijuana-323 dataset. However, linear models perform better in accuracy, precision, and recall. However, only GAT performs robust classification in the HCP datasets.
In our ablation experiments, we verified the contribution of the sub-models of the MixHop and LSTM sub-models in our overall model framework. First, we removed the high-order fusion of the MixHop model in the HOGANN and observed a substantial improvement in AUC, reaching 81.9% and 70.4% compared to the graph-based baseline. This indicates that the overall predictive capability of the model is limited when relying solely on the sequential LSTM. Additionally, it compared the BrainGB framework, specifically designed for the brain network, with different message-passing or pooling strategies. The experimental results from BrainGB indicate that utilizing only the one-hop neighborhood of local nodes achieves an AUC of 83.4% in the Marijuana-323 dataset but fails to perform adequately in the HCP dataset. Meanwhile, precision and recall are lower than the original GAT model, with values of 73.6% and 69.3%, respectively. Next, we eliminated the influence of LSTM in LOGAB. It was observed that incorporating MixHop into the model substantially enhanced the AUC for overall brain network learning, achieving 85.1% for marijuana-323 and 75.5% for the HCP dataset, with accuracies of 80.7% and 71.2%, respectively.
One of the advantages of HOGANN is its ability to integrate two sub-models: one dedicated to extracting temporal graph features and another focused on learning higher-order attention, ultimately combining these features for more effective messaging. Experimental results demonstrate its superiority over linear models and other graph-based methods, resulting in increased classification accuracy at 80.7%, precision at 81.0%, and recall at 80.7% compared to state-of-the-art models.
Next, we assessed the cross-validation performance of HOGANN by partitioning the complete dataset into five folds and conducting a 5-fold cross-validation. As depicted in Fig. 7, our HOGANN exhibits a notable AUC of 82.4% $\pm$ 0.029 in multigraph classification.
3.3 Classification based on varying length of fMRI time series
This section explores the impact of segmenting full-time series and subsequences into multigraphs with different window sizes $T^{\prime}$. Our training strategy employs four-time windows of length $T^{\prime}\in$0, 100, 300, 600} for training and testing. As depicted in Fig. 4, it is evident that the AUC, accuracy, F1, and precision exhibit substantial reductions as the number of segmented fMRI signals increases (segment $>$ 3) in the marijuana-323. However, GAT demonstrates a more stable performance when segmenting two or three times compared to GCN. HOGANN showcases increasingly robust classification performance in our experiments as the number of segments increases. Specifically, under the optimal condition of segmenting fMRI signals twice, HOGANN achieves an AUC of 82.7%, accuracy of 79.2%, F1 of 79.26%, and precision of 79.3%.
3.4 Ablation study
We conducted ablation studies on the sub-model effect to explore the impact of the loss function. Several optimization experiments were designed to compare the effectiveness of Eq. (19) on the overall model learning, as presented in Table 2. Subsequently, attention mechanism ablation experiments were conducted on the effect of multi-head attention, with each head independently attending to information from neighboring nodes’ latent space. Following that, we compared the classification effect of different $K$ nearest neighbor parameters on multigraph classification in two datasets.

### v5

@@ 2.2.4 Brain network reconstruction and weighted functional connectivity
where $L^{c}_{i}\in\mathbb{R}^{1\times d}$ can be expressed features maps for each ROI, we utilize the final output node embedding, denoted by $\textbf{L}^{c}=\left[L^{c}_{1},L^{c}_{2},...,L^{c}_{v}\right]$ for brain network reconstruction. This is accomplished through an inner product operation, as depicted in Eq. (17):
$\textbf{B}_{R}=\textbf{L}^{c}\cdot(\textbf{L}^{c})^{{T}},$ (17)
where $\textbf{B}_{R}\in\mathbb{R}^{v\times v}$ represents the global brain reconstruction matrix for ROIs. To enhance the structural consistency of ROIs across subjects in functional connectivity analysis, we employ element-wise multiplication with the reconstruction ROI matrix $\textbf{B}_{R}$. This operation is applied to each subject’s connectivity matrix $\textbf{A}^{(i)}(t)$ at time step $t$, where $i$ represents the subject index and $t$ denotes the temporal segment. This computation is shown in Eq. (18), and the subject-specific weighted functional connectivity matrices can be expressed as:
$\mathbf{W}_{fc}^{(i)}(t)=\mathbf{A}^{(i)}(t)\odot\sigma(\textbf{B}_{R}),$ (18)
where $\sigma(\cdot)$ is the sigmoid function, which normalizes $\textbf{B}_{R}$ to a range between 0 and 1, and $\odot$ denotes the element-wise product. Subsequently, we average these weighted functional connectivity matrices over time as follows:
$\mathbf{\bar{W}}_{fc}^{(i)}=\frac{1}{N}\sum_{t=1}^{N}\mathbf{W}_{fc}^{(i)}(t),$ (19)
where $\mathbf{\bar{W}}_{fc}^{(i)}$ represents the final average weighted functional connectivity (AWFC) matrix. The advantage of constructing an AWFC matrix is its ability to capture an overall weighted representation of the brain network. This representation enables further analysis of degree centrality (DC) and network community differences between distinct groups based on edge density analysis.
2.3 Objective function
@@
We can employ the output embeddings from Eq. (6) and Eq. (15) throughout the training phase. We then use Eq. (20) and Eq. (21) to execute a binary classification task, incorporating the final non-linear transformation $\sigma(\cdot)$ as follows:
$\hat{\textbf{Y}}_{c}=\sigma(\textbf{W}_{c}\textbf{H}^{c}+b_{c}),$ (20)
$\hat{\textbf{Y}}_{MixGAT}=\sigma(\textbf{W}_{a}\textbf{H}^{a}+b_{a}),$ (21)
@@
where $\textbf{W}_{c}$ and $\textbf{W}_{a}$ are the trained parameter weights from the two sub-models, while $b_{c}$ and $b_{a}$ represent their respective biases. Subsequently, our objective is to minimize the overall loss function derived from the losses of two sub-models. The first sub-model pertains to the MixHop attention loss, denoted as $\mathcal{L}_{MixGAT}$, while the second sub-model involves the combination fusion model loss denoted as $\mathcal{L}_{c}$. The combined loss can be expressed as:
@@
where Y represents the ground truth label, while $\hat{\textbf{Y}}_{MixGAT}$ and $\hat{\textbf{Y}}_{c}$ represent the class predictions for marijuana users and healthy controls, respectively. By combining $\mathcal{L}_{MixGAT}$ and $\mathcal{L}_{c}$, we can better leverage the fusion information in the total loss $\mathcal{L}$ not just sub-model loss. We will discuss the impact of loss on HOGANN overall prediction performance in more detail in Section 3.1.
Table 1: Summary statistics of demographics in Marijuana-323 and HCP
Cohort Dataset Age (Mean ± SD) Gender (Female/Male)
HC Marijuana-323 (n=128) 30.06 ± 10.47 73/55
HCP (n=991) 28.82 ± 3.72 567/424
LM Marijuana-323 (n=195) 27.23 ± 7.73 73/122
HCP (n=100) 28.45 ± 3.49 26/74
Significance Marijuana-323 $t=2.79^{\star}$ $\chi^{2}=11.20^{\star\star}$
HCP $t=0.95$ $\chi^{2}=34.42^{\star\star}$
[…]
2.4 Datasets
  • Marijuana-323: In this study, we extended the prior work of Kaustubh R. et al [34, 35, 36]. The Institutional Review Board (IRB) of this study was exempted by the Stevens Institute of Technology Review Board review committee. The research involved the collection of data from two datasets of fMRI images comprising 125 and 198 participants, recruited either from the non-medical marijuana-seeking community or from hospitalized patients. Participants were required to withdraw from marijuana for 72 hours and then asked to complete a cue-elicited craving task during an MRI scan. After the MRI scans, participants completed the marijuana craving questionnaire [37], marijuana withdrawal checklist, and marijuana problem investigation [38]. The data underwent reprocessing from 3T imaging fMRI data, and an average time series of 90 ROIs was calculated for each subject based on the Stanford atlas  [39]. In our final training and testing, there were 195 long-term marijuana users and 128 healthy control individuals.
  • HCP dataset: To assess the effectiveness of HOGANN, we evaluated an external dataset sourced from the human connectome project (HCP) S1200 [40], comprising 1096 rs-fMRI scans of young adults. The S1200 constitutes the major releases from 2012 to 2015 from Washington University in St. Louis, Missouri, United States. The HCP dataset encompasses 598 individuals who have used marijuana, with an average age of 28.76 ± 3.69, and 493 individuals who have never used marijuana, with an average age of 28.82 ± 3.72. This data preparation utilized each subject’s first session and excluded 5 rs-fMRIs containing fewer than 1200 frames. The cortical surfaces of each subject were partitioned into 22 brain regions serving as ROIs  [41], and blood oxygenation level-dependent (BOLD) time series signals were normalized using z-scores  [20].
@@ 3 Results
Table 1 displays summary statistics of the demographics for two datasets. In our experimental setup, the Marijuana-323 and HCP datasets were divided into training and testing sets in an 80/20 ratio. The Marijuana-323 dataset comprised 258 training subjects and 65 testing subjects, while the HCP dataset comprised 876 training subjects and 218 testing subjects.
Table 2: Evaluating classification performance with five-fold cross-validation. We computed the most competitive baseline for each method. We compared the second-best methods denoted by $\blacktriangle$" and calculated the improvement rate, denoted as "Improv. (%)". Here, $\star$" denotes statistical significance at the 0.05 level using a two-sample t-test.
Marijuana-323 (90 ROIs) HCP (22 ROIs)
Model AUC Acc. (%) Prec. (%) Rec. (%) AUC Acc. (%) Prec. (%) Rec. (%)
LR + L1 0.71$\pm$0.08 73.4$\pm$7.5 74.3$\pm$7.5 73.5$\pm$7.5 0.62$\pm$0.02 62.0$\pm$1.9 62.0$\pm$5.6 68.5$\pm$17.1
LR + L2 0.72$\pm$0.09 74.4$\pm$8.2 75.5$\pm$9.0 74.4$\pm$8.3 0.57$\pm$0.04 57.1$\pm$4.1 66.9$\pm$14.3 53.9$\pm$35.9
SVM + L1 0.68$\pm$0.08 69.7$\pm$7.5 73.6$\pm$9.5 73.1$\pm$9.0 0.64$\pm$0.01 63.2$\pm$0.8 64.1$\pm$1.2 64.1$\pm$1.2
SVM + L2 0.71$\pm$0.10 73.0$\pm$8.9 73.6$\pm$9.5 73.1$\pm$9.0 0.64$\pm$0.01 63.7$\pm$0.9 63.6$\pm$1.0 62.6$\pm$1.9
XGboost 0.65$\pm$0.01 67.6$\pm$8.3 66.6$\pm$14.4 67.6$\pm$8.3 0.65$\pm$0.11 67.6$\pm$8.3 66.6$\pm$14.4 67.6$\pm$8.3
GCN 0.77$\pm$0.01 68.4$\pm$4.5 75.7$\pm$8.4 74.8$\pm$11.8 0.66$\pm$0.06 68.0$\pm$2.6 69.7$\pm$2.4 74.0$\pm$3.2
GAT 0.79$\pm$0.04 72.1$\pm$3.8 75.3$\pm$6.2 81.7$\pm$9.4 0.67$\pm$0.02 72.7$\pm$1.6 73.1$\pm$0.9 75.1$\pm$4.9
GraphSAGE 0.77$\pm$0.00 64.6$\pm$5.3 73.6$\pm$5.6 68.2$\pm$21.2 0.74$\pm$0.01 67.5$\pm$2.3 68.4$\pm$1.4 67.4$\pm$8.5
BrainGB 0.72$\pm$0.04 66.0$\pm$2.2 65.5$\pm$2.7 64.1$\pm$2.8 0.62$\pm$0.01 59.1$\pm$0.7 59.0$\pm$1.2 57.4$\pm$0.9
BrainUSL 0.64$\pm$0.19 62.7$\pm$9.6 71.3$\pm$9.3 73.3$\pm$30.0 0.48$\pm$0.26 73.2$\pm$12.0 73.0$\pm$9.4 79.5$\pm$16.0
Graph-Transformer 0.75$\pm$0.04 70.9$\pm$3.5 75.1$\pm$5.4 77.9$\pm$2.4 0.76$\pm$0.06 72.4$\pm$2.6 72.4$\pm$4.8 77.8$\pm$5.2
BrainGSL 0.80$\pm$0.02 74.3$\pm$2.8 77.5$\pm$2.2 80.7$\pm$6.5 0.77$\pm$0.06 74.5$\pm$9.6 74.3$\pm$6.0 81.3$\pm$18.3
HOGANN (w/o GAT-LSTM) 0.80$\pm$0.02 74.7$\pm$2.8 77.3$\pm$3.1 82.5$\pm$7.0 0.74$\pm$0.06 73.2$\pm$5.7 74.5$\pm$5.9 78.6$\pm$8.0
HOGANN (w/o MixHop) 0.78$\pm$0.03 70.1$\pm$4.4 77.3$\pm$3.0 83.8$\pm$4.5 0.79$\pm$0.04 72.7$\pm$5.3 73.6$\pm$5.0 79.0$\pm$5.3
HOGANN 0.86$\pm$0.03 79.2$\pm$3.6 79.7$\pm$4.0 87.9$\pm$1.3 0.85$\pm$0.04 78.5$\pm$2.2 81.2$\pm$3.4 83.0$\pm$3.4
Improv. (%) 7.5% 6.5% 2.8% 8.9% 10.3% 5.3% 9.5% 2.0%
Significance $\star$ $\star$ $\star$ $\star$ $\star$
3.0.1 Classical methods and graph-level classification
The experiment results of the proposed HOGANN, as shown in Table 2, illustrate the model performance across different approaches in the marijuana-323 (90 ROIs) and HCP (22 ROIs) datasets. We first used ML methods as the baseline, our experiments reveal that among the ML methods: LR + L1, LR + L2, SVM + L1, SVM + L2 and XGboost. Particularly, LR + L2 achieved the highest classification accuracy at 74.4% in Marijuana-323, as depicted in Table 2. However, the same ML methods did not yield superior results in the HCP dataset. These classification results across the two datasets indicate the presence of nonlinear structures among brain regions in the fMRI scans of LM users, which ML methods fail to discern based on functional connectivity features. Finally, to evaluate the graph topology structure of fMRI data, we reasonably compare graph-based deep learning with our proposed HOGANN, such as GCN, GAT, GraphSAGE, and BrainGB as baseline models. Table 2 shows that GAT achieve better area under the curve (AUC) scores than ML methods in the marijuana-323 dataset. The experimental results from GAT indicate that utilizing only the one-hop neighborhood of local nodes achieves an AUC of 0.79 in the Marijuana-323 dataset but fails to perform adequately in the HCP dataset (AUC=0.67). Meanwhile, BrainGB and GraphSAGE with mini-batch node feature sampling methods also fail to outperform GAT, which has an attention mechanism, in terms of performance.
3.0.2 Compare with SOTA models
Furthermore, we conduct a fair comparison between our HOGANN and more complex graph deep learning architectures, including Graph-Transformer, BrainUSL, and BrainGSL) on two datasets, using a fixed learning rate of 0.01. Based on the experimental results, BrainUSL, an unsupervised framework, exhibits inferior prediction performance on the Marijuana-323 and HCP datasets compared to the ML method when evaluated using AUC scores. We can reasonably explain that unsupervised models are prone to generalization failures when applied across different datasets, leading to poor model performance. Furthermore, we compared the second-best model, BrainGSL, with self-supervised learning to our HOGANN. The comparison resulted in significant improvements: an 7.5% increase in AUC on the Marijuana-323 dataset and a 10.3% increase on the HCP dataset. Additionally, on the HCP dataset, we observed improvements of 5.3% in accuracy and 9.5% in precision. However, these models typically concentrate solely on the topological information extracted from the entire fMRI time series and do not consider the dynamic topological structure in different time intervals. To address these limitations and validate our approach, we present comprehensive ablation studies in the following section 3.1.
Figure 3: The impact of KL divergence ablation studies on mixed embeddings $\textbf{H}^{a}$ and $\textbf{H}^{\prime}$: (Left) Comparing KL divergence with and w/o MixHop. (Right) Comparing KL divergence with weighted loss versus unweighted loss.
Figure 4: The ablation analysis of HOGANN classification in LM and HC classification. Panel (A) demonstrates the performance changes with multiple attention heads, while Panel (B) illustrates the classification performance changes when the neighborhood K in of the HOGANN is varied. Panel (C) comparing different models based on their classification performance on the marijuana-323 with varying multi-segment time window sizes.
Figure 5: A comparison of machine learning methods with our HOGANN in identifying the optimal community clustering. In Panel (A) and (B), we compare community detection in weighted functional connectivity matrices using machine learning and our HOGANN, where we map the most significant four communities of the LM group to their corresponding brain regions of the connectivity network. Based on Panel (C) for the LM group, we depict four communities corresponding to the colored boxes, illustrating the connectivity of the most active brain regions in Panel (D).
Table 3: The ablation experiment in loss comparison
Marijuana-323 (90 ROIs) HCP (22 ROIs)
Strategies AUC Accuracy (%) Precision (%) Recall (%) AUC ACC (%) Precision (%) Recall (%)
$\mathcal{L}_{c}$ 0.79 76.9 76.7 76.9 0.66 63.9 63.9 63.9
$\mathcal{L}_{MixGAT}$ 0.79 73.9 73.7 73.9 0.62 68.0 68.3 68.0
$\alpha\cdot\mathcal{L}_{MixGAT}+(1-\alpha)\cdot\mathcal{L}_{c}$ 0.79 75.4 75.4 75.4 0.59 62.5 62.5 62.5
$\mathcal{L}_{c}+\mathcal{L}_{MixGAT}$ 0.84 78.5 78.3 78.5 0.70 69.1 69.4 69.1
3.1 Ablation studies
In our ablation experiments, we observed that utilizing MixHop or GAT-LSTM independently yielded superior classification performance improvements across various metrics in the LM and HC groups compared to most graph-based model baselines (such as GCN, GAT, GraphSAGE, and BrainGB). However, we noted that the simultaneous integration of higher-order attention mixing and dynamic sequential graph approach enables the preservation of high-order attention mechanisms in overall brain network learning for more effective messaging. Our research conducts the following ablation experiments on HOGANN’s hyperparameters, loss function, and the number of segments in the fMRI time series: